OpenStudy (anonymous):

modular arithmetic operations 8 ∙ 8 (mod 6) I came up with 4 by using a calculator but if someone could briefly explain what mod is or why to use it that would be awesome!

OpenStudy (nincompoop):

don't use a calculator

OpenStudy (anonymous):

im trying really hard to understand, is this right? 8*8(mod6) 64/6=10.66 so the answer is 6?

OpenStudy (anonymous):

8*8(mod6) = 64(mod 6) so we can rewrite it as: 64 - 6k = r where k\[k \in Z\] and r is the remainder so \[\lfloor \frac{ 64 }{ 6 } \rfloor\] = 10 so let k = 10 so 10*6 = 60 so r = 64-60 =4

OpenStudy (amistre64):

mod 6 just means that we only have 6 numbers to count with, usually the set: 0,1,2,3,4,5 if we want to count beyond 5, we have to circle about back to the start 0, 1, 2, 3, 4, 5 6, 7, 8, 9, 10, 11 if we want to count beyond 11, we have to circle about back to the start 0, 1, 2, 3, 4, 5 6, 7, 8, 9, 10, 11 12, 13, 14, 15, 16, 17 this cyclical movement defines equivalence classes; for any number in a given column, it is equivalent to (or congruent to) the number presented at the top of the column. notice that the number 8 is in column 2; so we say that 8 is congruent to 2, mod6 8 = 2 mod6 8*8 = 2*2 mod6